# Decentralizing Test-time Adaptation under Heterogeneous Data Streams

### v1

@@ C Continual Setting Evaluation
Although our method is specifically designed for mixed domain scenarios, we also evaluated its performance under the conventional continual test-time adaptation setting to assess its robustness in different contexts. In this setting, the model adapts online to a sequence of test domains without explicit knowledge of domain shifts, with only one distribution shift occurring at a time and not reappearing. Without adjusting any parameters, our method demonstrated competitive performance compared to current state-of-the-art approaches.
| CLUSTER_NUM | 2 | 4 | 8 | 16 |
| CIFAR10-C | 23.0 | 22.9 | 23.2 | 24.7 |
| CIFAR100-C | 34.8 | 34.7 | 34.7 | 35.6 |
| IN-C (BN) | 68.6 | 67.2 | 67.1 | 70.5 |
| IN-C (LN) | 50.3 | 48.8 | 49.9 | 50.0 |
| KMEANS_SIZE | 256 | 512 | 1024 | 2048 |
| CIFAR10-C | 23.0 | 22.9 | 23.0 | 22.9 |
| CIFAR100-C | 34.6 | 34.7 | 34.8 | 34.8 |
| IN-C (BN) | 69.0 | 67.2 | 67.6 | 67.0 |
| IN-C (LN) | 49.0 | 48.8 | 48.7 | 48.8 |
| COMM_INTERVAL | 1 | 10 | 100 | 1000 |
| CIFAR10-C | 22.6 | 22.9 | 22.6 | 22.0 |
| CIFAR100-C | 34.7 | 34.7 | 34.9 | 43.2 |
| IN-C (BN) | 67.1 | 67.2 | 67.2 | 67.4 |
| IN-C (LN) | 48.4 | 48.8 | 48.8 | 48.7 |
Notably, while UnMix-TNS effectively addresses non-i.i.d. issues (dependent sampling at the class level), it is less effective under i.i.d. conditions. Our results suggest that the proposed FreDA not only excels in its intended mixed domain scenarios but also generalizes effectively to standard continual adaptation tasks, providing a robust solution across various distributional challenges.
D Parameter Study
In this section, we study the parameter choice of CLUSTER_NUM, KMEANS_SIZE and COMM_INTERVAL (refer to Algorithm 1 for detailed definitions). Results are reported in Table 6.
As we adjust the KMEANS_SIZE parameter from 256 to 2048, there is a remarkably consistent performance on different datasets, indicating that our method’s capability to generalize across various sizes.
The variation in CLUSTER_NUM across our datasets underscores the nuanced balance required in selecting the optimal branch number for domain adaptation. Utilizing just two clusters already yields relatively good results, suggesting that a minimal decentralization can be effective. However, as the number of clusters increases from 2 to 16, we observe a decline in performance on CIFAR100-C and a more pronounced deterioration on ImageNet-C, with the optimal performance achieved at a CLUSTER_NUM of 4. This trend underscores the delicate trade-off between model complexity and the risk of overfitting: employing too large a cluster size can lead to a model overly tailored to the training data, impairing its generalization capabilities.
For the sensitivity analysis of COMM_INTERVAL, we observe that our method is generally robust to changes in the communication interval across all datasets. However, the impact of communication frequency varies significantly among different datasets. For simpler datasets like CIFAR10, minimal communication, exemplified by an interval of , yields the best results. This could be attributed to the model’s high accuracy, enabling positive feedback loops even within isolated branches. Conversely, for more complex datasets, more frequent communication, with intervals as low as , appears beneficial. This frequent updating may help prevent model degradation over time, especially in scenarios where the data complexity could lead to significant divergences in learning pathways among distributed model components.
E Adaptation Scenarios
@@
Mixed Domains: In this scenario, the model processes a long sequence of test samples where each sample is independently drawn from a randomly selected target domain and a randomly selected class among classes . The sequence is represented as:
@@
Mixed Domains & Dependent Sampling: This scenario, extending the mixed distribution framework, introduces time-correlated sequential data emergence from the same class, featuring both covariate and label shifts. While the domain indices are independently and randomly selected for each sample as before, the class labels are correlated over sequences of variable lengths. Specifically, the samples tend to come from the same class for a variable number of consecutive samples before potentially switching to a different class. The sequence can be represented as:
@@
are independently and randomly selected target domain index.
refers to the size of the -th data segment in the sequence where .
denotes the class of samples within the -th data segment.
F Dataset Visualization
@@
G Limitation and Future Work
While FreDA addresses a critical challenge in handling heterogeneous data streams, providing a solid pipeline for this issue, there are still avenues for further enhancement.
Our current aggregation approach, which averages models based on cluster counts, has been effective in solving the problem at hand. However, exploring alternative strategies—such as weighting models by the divergence between clusters—might lead to incremental improvements. Additionally, refining the sample selection process from a original sample-level focus to a more granular patch-level could extend FreDA’s applicability to tasks such as segmentation, thereby further enhancing its versatility in real-world scenarios.

### v2

@@ VII-C3 Performance under continual settings
Although our method is specifically designed for mixed domain scenarios, we also evaluated its performance under the conventional continual test-time adaptation [5, 3] setting to assess its robustness in different contexts. In this setting, the model adapts online to a sequence of test domains without explicit knowledge of domain shifts, with only one distribution shift occurring at a time and not reappearing. Without adjusting any parameters, our method demonstrated competitive performance compared to current state-of-the-art approaches. Notably, while UnMix-TNS effectively addresses non-i.i.d. issues (dependent sampling at the class level), it is less effective under i.i.d. conditions. Our results suggest that the proposed FreDA not only excels in its intended mixed domain scenarios but also generalizes effectively to standard continual adaptation tasks, providing a robust solution across various distributional challenges.
VIII Conclusion and Future Work
This paper advances Test-Time Adaptation (TTA) by tackling the real-world complexities of heterogeneous data streams. Our decentralized approach precisely manages diverse data shifts, improving model adaptation in varied settings. By integrating Fourier-based augmentation, we expand the range of confident samples for each distribution shift, further boosting model performance. The experimental results underscore the efficacy of FreDA, highlighting its potential to influence the field and guide future research in adapting to dynamic and diverse data shifts. While FreDA addresses a critical challenge in handling heterogeneous data streams, providing a solid pipeline for this issue, there are still avenues for further enhancement. Our current aggregation approach, which averages models based on cluster counts, has been effective in solving the problem at hand. However, exploring alternative strategies – such as weighting models by the divergence between clusters – might lead to incremental improvements. Additionally, refining the sample selection process from a original sample-level focus to a more granular patch-level could extend FreDA’s applicability to tasks such as segmentation, further enhancing its versatility in real-world scenarios.
-A Adaptation Scenarios
-A1 Mixed Domains
@@ VIII Conclusion and Future Work
: In this scenario, the model processes a long sequence of test samples where each sample is independently drawn from a randomly selected target domain and a randomly selected class among classes . The sequence is represented as:
@@
-A2 Continual Domain Adaptation
: In this setting, the model adapts online to a sequentially presented series of test domains, where each domain shift occurs only once and does not reappear. The sequence progresses through distinct target domains in a fixed order , with samples from each domain appearing contiguously. The class labels within each domain are independently and randomly selected. The sequence is structured as:
@@
denotes the -th target domain in the fixed sequence, with .
is the randomly selected class label for the -th sample, independent of domain transitions.
-B Dataset Visualization
@@
-C Relation to Frequency Domain Learning
Recent advances highlight frequency-based techniques as powerful tools for domain transfer. In domain generalization, frequency analysis has revealed critical insights into model robustness and learning dynamics [46, 47, 48, 49, 50, 51, 52, 53]. For domain adaptation, interpolating image amplitude spectra across styles has proven effective in reducing domain gaps and preventing overfitting to low-level statistics [54, 55, 56, 57]. Motivated by these advancements, we discover that frequency information inherently captures domain characteristics, making it a valuable medium for decoupling mixed target domains – an aspect largely unexplored in prior work. On this basis, we propose FreDA, the first framework addressing TTA under mixed domain shifts in Fourier space, with a decentralized adaptation and perturbation mechanism.
-D Relation to Multi-Target Domain Adaptation
TTA under mixed distribution shifts shares similarities with multi-target unsupervised domain adaptation (MT-UDA) [11, 12, 58, 59], yet diverges critically in complexity: while MT-UDA assumes static, predefined target domains and leverages labeled source data for explicit domain alignment, TTA operates with no access to source data and must adapt to dynamic, unpredicTABLE target streams in an online manner. This eliminates direct source-target discrepancy computation and demand robust incremental adaptation rather than offline multi-domain optimization. These differences highlight the need for novel methodologies tailored specifically to TTA, going beyond the solutions developed for MT-UDA.
-E Relation to Decentralized, Federated, Distributed Learning
This work also intersects with decentralized, federated, and distributed learning due to splitting data batches into disjoint subsets and applying decentralized model adaptation. First, while decentralized learning focuses on non-i.i.d. data that is naturally distributed across multiple nodes [60], our approach starts with a centralized batch of target samples. By proactively splitting this data into disjoint subsets, we expose latent non-i.i.d. characteristics, enabling the effective use of decentralized learning techniques. Second, federated learning considers data privacy and model collaborations within decentralized learning [61]. In our case, as target samples are mixed in a batch, data privacy is not a concern. However, similar to federated learning, our approach also involves weight aggregation over subnetworks to enhance their base models before next batch adaptation. Third, distributed learning aims to improve training efficiency on large-scale datasets by partitioning data for synchronized training [62]. In contrast, our method operates in a real-time fine-tuning context with limited data at one time, hence scalability is less of a concern.